OpenStudy (anonymous):

Monisha decides to join a gym for a minimum of one month to train for a triathlon. The gym charges a joining fee of $100 and a monthly fee of $37. If x represents the number of months that Monisha is a member of the gym, the equation below can be used to determine C, her total membership fee for that duration of time. 100 + 37x = C Monisha has allocated a maximum of $285 to spend on her gym membership. Which number line shows the possible number of months that Monisha can be a member of the gym? Answer

OpenStudy (4n1m0s1ty):

Can you post the answers?

OpenStudy (anonymous):

no it is 4 graphs

OpenStudy (4n1m0s1ty):

Well its kind of hard for me to help, when I don't know what the number lines look like

OpenStudy (anonymous):

If C = {1, 2, 3, 4, 5, 6, 7, 8, 9, 10} and D = {numbers not divisible by 2 from 1 to 10}, what is CD? Answer {1, 2, 3, 4, 5, 6, 7, 8, 9, 10} {1, 3, 5, 7, 9} { } {2, 4, 6, 8, 10}

OpenStudy (4n1m0s1ty):

2nd one

OpenStudy (anonymous):

your awsome Let A = {t-shirt, tank} and let B = {shorts, pants}. Find A X B. Answer {(t-shirt, short), (tank, pants)} {t-shirt, tank, shorts, pants} {(t-shirt, shorts), (t-shirt, pants), (tank, shorts), (tank, pants)} { }

OpenStudy (4n1m0s1ty):

3rd one i think. Cartesian Product? havent done this stuff in a while

OpenStudy (anonymous):

thanks Which relation does not represent a function? Answer (-3, 6), (-1, 1), (0, 1), (3, 7) (-5, 3), (-3, 1), (0, -4), (3, -6) (-4, 2), (1, 5), (-2, 0), (1, 6) (2, 0), (4, 6), (6, 8), (10, 6)

OpenStudy (4n1m0s1ty):

Functions means that there is an exactly 1 x for each y. 3rd one

OpenStudy (4n1m0s1ty):

you have (1,5) and (1,6)

OpenStudy (anonymous):

makes since What is the slope between the two points (4, 13) and (1, 4)? Answer 3 -1/3 1/3 -3

OpenStudy (4n1m0s1ty):

slope = change in y over change in x (13 - 4)/(4 -1) = 9/3 = 3

OpenStudy (anonymous):

What is the slope and y-intercept of the graph of 3x + 2y = 12? Answer slope =-2/3 y-intercept = (0, 6) slope = 6 y-intercept = (0,-3/2 ) slope = -3; y-intercept = (0, 12) slope = -3/2 y-intercept = (0, 6)

OpenStudy (4n1m0s1ty):

Just rearrange to y = mx + b form 2y = -3x +12 y = (-3/2)x + 6 (last one)

OpenStudy (anonymous):

your great What is the equation of the line that passes through the point (3, 2) and has a slope of 4? Answer y = 4x - 5 y = 4x - 1 y = 4x - 10 y = 4x + 1

OpenStudy (4n1m0s1ty):

y-2 = 4(x-3) --> y - 2 = 4x - 12 --> y = 4x -10 Medals plz

OpenStudy (anonymous):

What is the equation of the line that passes through the points (-2, 1) and (1, 10)? Answer 3x - y = 5 3x - y = -7 x + 3y = -5 3x - y = -5

OpenStudy (anonymous):

you can do that GOLD

OpenStudy (4n1m0s1ty):

Just did guess and check to lazy for algebra right now... 2nd one

OpenStudy (4n1m0s1ty):

you in high school btw?

OpenStudy (anonymous):

yes Select the equations that are parallel and perpendicular to y = x + 5 and that pass through the point (-2, -1). Answer parallel: y = x - 1 perpendicular: y = -x + 1 parallel: y = -x - 1 perpendicular: y = x + 2 parallel: y = x + 1 perpendicular: y = -x - 3 parallel: y = 2x - 2 perpendicular: y = -2x - 1

OpenStudy (anonymous):

and am horrible at it

OpenStudy (4n1m0s1ty):

3rd one i think. Running on 2 hours of sleep, after cramming for one of my finals

OpenStudy (4n1m0s1ty):

so don't take my word on that one

OpenStudy (anonymous):

i belive you....your smart Choose the correct description of the graph of the compound inequality 3x < -9 or x + 2 10 Answer A number line with a closed circle on -3, an open circle on 8, and shading in between. A number line with a closed circle on -3, shading to the left, and an open circle on 8, shading to the right. A number line with an open circle on -3, a closed circle on 8, and shading in between. A number line with an open circle on -3, shading to the left, and a closed circle on 8, shading to the right.

OpenStudy (anonymous):

are you in collage?

OpenStudy (4n1m0s1ty):

yeah 2nd year whats the 2nd part suppose to say x + 2 10 ?

OpenStudy (anonymous):

>=

OpenStudy (4n1m0s1ty):

last one

OpenStudy (4n1m0s1ty):

i should be working on a take home final right but im too tired :(

OpenStudy (anonymous):

i can help :) Choose the correct description of the graph of the compound inequality x - 2 > -10 and 2x >= 12 Answer A number line with a closed circle on -8, shading to the left, and an open circle on 6, shading to the right. A number line with a closed circle on -8, an open circle on 6, and shading in between. A number line with an open circle on -8, a closed circle on 6, and shading in between. A number line with an open circle on -8, shading to the left, and a closed circle on 6, shading to the right.

OpenStudy (anonymous):

no it is <=

OpenStudy (anonymous):

x - 2 > -10 and 2x <= 12

OpenStudy (4n1m0s1ty):

3rd one

OpenStudy (anonymous):

your the best Determine the type of boundary line and shading for the graph of the inequality 4x + y > -6. Answer Dashed line with shading on the side that does not include the origin. Dashed line with shading on the side that includes the origin. Solid line with shading on the side that does not include the origin. Solid line with shading on the side that includes the origin.

OpenStudy (4n1m0s1ty):

hate these. lets skip this one. how many more do you have left?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

one after this one and it is a hard one...for me anyways Connie needs to burn at least 400 calories at the gym today. She can burn 20 calories per minute on the treadmill and 12 calories per minute on the elliptical trainer. Which inequality below can be used to find the possible combinations of time she could spend on each machine and still meet her goal of burning at least 400 calories? Answer 20t - 12e 400 20t + 12e 400 20t + 12e 400 20t - 12e 400

OpenStudy (4n1m0s1ty):

20t + 12e >= 400

OpenStudy (4n1m0s1ty):

or just >

OpenStudy (anonymous):

last one its a three parter Assume a 115 pound person burns 36 calories after 12 minutes of walking and 54 calories after 18 minutes of walking. Part 1 : Use calories as the y-coordinate and minutes as the x-coordinate. After writing two ordered pairs, find the slope. What does the slope represent in terms of the information provided? Part 2 : Write an equation, in slope intercept form, to represent this data. Part 3 : How many calories will this person burn after 35 minutes of walking? Using complete sentences, explain how the equation, slope, or graph can help to predict calories burned.

OpenStudy (4n1m0s1ty):

Part 1: (12, 36) and (18, 54); slope = (54 -36)/(18-12) = 18/6 = 3 Part 2: y - 36 = 3(x - 12) --> y = 3x - 36 + 36 --> y = 3x Part 3: Just plug in 35 for x in part 2. I think you can explain how the model works.

OpenStudy (anonymous):

you good at history :)

OpenStudy (4n1m0s1ty):

depends, what part?

OpenStudy (4n1m0s1ty):

I'm an engineer not a history major so I have my limits

OpenStudy (anonymous):

oh with the math i just wanted to say I SO PASSED

OpenStudy (4n1m0s1ty):

was that a homework you had?

OpenStudy (anonymous):

more of a study exam

OpenStudy (anonymous):

slaves, civil war stuff like that

OpenStudy (4n1m0s1ty):

yeah not too good at that part in particular. Might want to try the history section, though I doubt anyones on right now.

OpenStudy (anonymous):

ok thanks for all of your help so what do i do now?

OpenStudy (4n1m0s1ty):

Hit the blue bar on top and click find mor subjects at bottom, then join history study group

OpenStudy (anonymous):

ok so am i suppose to rate you or something?

OpenStudy (4n1m0s1ty):

You can give me a medal. just hit best answer on any of my replies

OpenStudy (anonymous):

ok done thanks

OpenStudy (anonymous):

ok well cya around time to go to history

OpenStudy (4n1m0s1ty):

cya